OpenStudy (anonymous):

Roberta is wrapping gift boxes. Each gift box is a rectangular prism with a square base of side 8 inches. The difference between the heights of the two gift boxes is 9 inches. If the larger box has a surface area of 544 square inches, what is the difference in the surface area of the two boxes? 576 square inches 648 square inches 324 square inches 288 square inches

sam (.sam.):

 Area of larger box, 2(8)(8)+4(x+9)(8)=544 x=4 -------------------------- Area of smaller box, 2(8)(8)+4(8x)=256 Difference in area = 544-256=288